# L2M-Reg: Building-level Uncertainty-aware Registration of Outdoor LiDAR Point Clouds and Semantic 3D City Models

### v2

@@ Abstract
Accurate registration between LiDAR (Light Detection and Ranging) point clouds and semantic 3D city models is a fundamental topic in urban digital twinning and a prerequisite for downstream tasks, such as digital construction, change detection and model refinement. However, achieving accurate LiDAR-to-Model registration at individual building level remains challenging, particularly due to the generalization uncertainty in semantic 3D city models at the Level of Detail 2 (LoD2). This paper addresses this gap by proposing L2M-Reg, a plane-based fine registration method that explicitly accounts for model uncertainty. L2M-Reg consists of three key steps: establishing reliable plane correspondence, building a pseudo-plane-constrained Gauss–Helmert model, and adaptively estimating vertical translation. Experiments on three real-world datasets demonstrate that L2M-Reg is both more accurate and computationally efficient than existing ICP-based and plane-based methods. Overall, L2M-Reg provides a novel building-level solution regarding LiDAR-to-Model registration when model uncertainty is present.
@@ 1 Introduction
LiDAR (Light Detection and Ranging) point clouds and semantic 3D city models are two widely used digital representations that play important roles in urban digital twinning (Jeddoub et al., 2023; Ketzler et al., 2020). However, as heterogeneous data types, they differ in many aspects such as data format, acquisition methods, geometric features, and accuracy. Specifically, semantic 3D city models primarily contain generalized structural information and serve as simplified and abstract representations of physical buildings, whereas LiDAR point clouds accurately capture detailed geometric features, providing high-precision representations (Kada and McKinley, 2009; Xu and Stilla, 2021). Accurate and efficient registration between them constitutes a fundamental research topic, as well as an essential prerequisite for numerous downstream applications such as digital construction, change detection, model updating and enrichment, model reconstruction and texturing (Liu et al., 2024; Shao et al., 2024; Zhu et al., 2024; Wysocki et al., 2021; Kulmer et al., 2025).
Internationally, CityGML (City Geography Markup Language), as a standard developed by the Open Geospatial Consortium (OGC), is widely adopted for representing and managing semantic 3D city models (Kolbe et al., 2021; Gröger and Plümer, 2012). CityGML supports the modeling of urban objects by incorporating their 3D geometry, appearance, topology, and semantic information across four distinct Levels of Details (LoD) (Kolbe et al., 2021). Benefiting from the widespread adoption of CityGML, accurately georeferenced LoD2 models are publicly available in many countries and regions (Wysocki et al., 2024). These models are typically maintained by governmental agencies or professional institutions to ensure consistent quality and reliability. Consequently, in most LiDAR-to-Model registration tasks, existing georeferenced LoD2 models are utilized as reference data, to which up-to-date LiDAR point clouds are aligned.
When discussing LoD2 building models used for accurate registration, a particularly critical issue is their inherent uncertainty. The encountered uncertainty stems from the current model generation process, which typically relies on 2D building footprints from the cadastral registry. As illustrated in Figure 1, the building footprint is defined by the outermost structural elements.
@@
As seen in Figure 1, buildings typically consist of various structural elements from the foundation to the roof, such as wall plinth (in red), exterior facade (in blue), and overhanging roof (in orange). Cadastral surveying regulations define which of these elements should be included in the model. Typically, a horizontal offset exists between a building’s plinth and its upper facade. Depending on the architectural style, this offset can range from several centimeters to decimeters. Since LoD2 building models are generally reconstructed by extruding building footprints to prototypical roof shapes from Airborne Laser Scanning (ALS) data (Roschlaub and Batscheider, 2016), the modeled wall surfaces actually only align with the building’s plinths, not facades. Consequently, the building plinths that closely correspond to footprint should be more suited for establishing correspondences with modeled wall surfaces.
Previous studies on LiDAR-to-Model registration have commonly assumed that the LoD2 models served as reference data are error-free, thus neglecting this inherent uncertainty. Although this assumption is basically accepted, it will become problematic in high-precision application scenarios. For example, when the captured LiDAR point clouds need to be accurately georeferenced based on existing LoD2 models, this horizontal offset between the plinth and exterior facade cannot be disregarded. Recent research has acknowledged similar issues about model uncertainty, focusing primarily on visual and conceptual representations from architectural or modeling perspectives (Landes et al., 2019; Potter et al., 2012; Zou and Sester, 2022). However, most LiDAR-to-Model registration studies have overlooked this uncertainty issue, treating it as a conventional point cloud registration task. As a result, a disconnection arises between building modeling and LiDAR-to-Model registration research.
Since urban dwellings form the backbone of any city, most existing semantic 3D city models are composed of buildings (Biljecki et al., 2015). Compared to city-level LiDAR-to-Model registration, accurate and efficient building-level solution offers the potential to refine city models based on specific needs, thereby avoiding extensive data acquisition and substantially reducing costs. However, this transition introduces several challenges:
@@
    Limited availability of geometric features. Individual buildings or partial building segments typically contain fewer geometric primitives—such as facade planes—with limited diversity in orientation and quantity. This constraint makes feature extraction and correspondence establishment more difficult.
@@
These challenges underscore a clear research gap in achieving accurate building-level LiDAR-to-Model registration under model uncertainty. In this context, model uncertainty is defined as the inconsistency between the final model, produced through a predefined generation process, and the actual building object, specifically in terms of geometric and structural representation. To address this, a tailored solution named L2M-Reg is proposed. To the best of our knowledge, this is the first comprehensive study to explicitly account for the inherent uncertainty of reference LoD2 models in building-level registration. The main contributions are summarized as follows:
@@
    A plane-based LiDAR-to-Model fine registration method, L2M-Reg, tailored for individual buildings is proposed. It integrates reliable plane correspondence establishment, pseudo-plane-constraint Gauss–Helmert Model (GHM), and adaptable vertical translation estimation. By explicitly considering the uncertainty in LoD2 models, L2M-Reg achieves superior performance on three real-world datasets.
@@ 2.1.1 Registration Directly Using LoD2 Models
The registration between LiDAR point clouds and their corresponding LoD2 models essentially involves feature extraction and correspondence matching across heterogeneous data. A straightforward strategy to address this issue is to directly extract geometric features, such as points, lines, and planes, from the models and subsequently match these features with corresponding ones derived from the point clouds (Bueno et al., 2018; Sheik et al., 2022a, b; Monasse et al., 2023; Bosché, 2012; Lucks et al., 2021; Goebbels et al., 2019). The transformation parameters can then be computed based on these matched features. In particular, most man-made structures include numerous planar surfaces, which naturally facilitates the extraction and utilization of planar features (Sheik et al., 2022a; Qiao and Butt, 2023).
@@ 2.1.2 Registration Using Converted Point Clouds from LoD2 Models
ICP-based methods are extensively employed to register point clouds. Traditional ICP algorithms iteratively align corresponding points but often exhibit slow convergence and sensitivity to initial values and outliers (Besl and McKay, 1992). Thus, many enhanced variants have been developed to address these limitations. For instance, Point-to-Line ICP (Censi, 2008) and Point-to-Plane ICP (Low, 2004; Rusinkiewicz and Levoy, 2001) utilize local linear and planar features, respectively, thereby accelerating convergence and enhancing registration accuracy. Generalized ICP (GICP) further refines registration precision by statistically optimizing correspondences through probabilistic modeling of local surface geometry (Segal et al., 2009). Trimmed ICP (TriICP) effectively reduces the influence of outliers by selectively excluding a certain percentage of correspondences, thereby enhancing robustness in noisy scenarios (Chetverikov et al., 2002). More recently, a robust method based on generated planar patches and adaptive distance thresholds was proposed, which has significantly reduced the influence of surface changes on registration accuracy (Yang and Schwieger, 2023; Yang and Holst, 2025). Collectively, these ICP variants have progressively addressed the limitations in standard ICP, significantly improving the accuracy, efficiency, and practical applicability of point cloud registration tasks.
Compared with the aforementioned ICP-based methods, geometric feature-based methods utilize points, lines, curves, and planes extracted from point clouds to establish correspondences for transformation estimation. These methods are generally less sensitive to variations of initial alignments compared to ICP-based methods. In the building scenarios discussed in this paper, planar features offer notable advantages over other geometric primitives, such as point or line features, and have thus been commonly utilized in point cloud registration tasks (Sheik et al., 2022a; Bosché, 2012; Wujanz et al., 2018; Chen et al., 2019; Xu et al., 2025). This preference arises primarily for two reasons. First, man-made structures typically contain abundant planar elements, naturally facilitating the extraction and use of planar features. Second, planar features demonstrate higher resilience against surface noise and outliers.
@@ 2.2 Inherent Uncertainty of Models
Compared to the rapid growth in the creation and application of models, research on their uncertainty has lagged behind and not received equivalent attention (Zou and Sester, 2022). Initially, researchers from architecture and archaeology took the lead in visualizing uncertainty associated with models of historical buildings (Zuk et al., 2005; Kensek, 2007). Subsequently, approaches to quantifying uncertainty in models have continued to evolve, with increasingly comprehensive methods proposed, along with the introduction of the concept of Level of Uncertainty (LoU), which systematically characterizes uncertainty variations among different components of building models (Landes et al., 2019; Potter et al., 2012).
With growing demand for high-precision applications relying on semantic 3D city models such as navigation (Kulmer et al., 2025), trajectory estimation (Lucks et al., 2021), and change detection (Meyer et al., 2022), an increasing number of researchers have also begun to investigate the uncertainty inherent in these models (Foschi et al., 2024). For instance, probabilistic methods have been developed for quantifying uncertainty, primarily focusing on semantic labeling in 2D maps or building models (Di et al., 2022; Feng et al., 2021; Paz et al., 2020). Additionally, other researchers have specifically examined the uncertainty associated with facades and windows in semantic 3D city models, quantifying aspects such as their positions and orientations (Zou and Sester, 2022).
Collectively, uncertainty in semantic 3D city models has become increasingly important in high-precision applications, as uncertainty inherent in the model inevitably propagates into downstream use cases. Neglecting these uncertainties or treating the model as an error-free reference is likely to introduce systematic inaccuracies and errors in subsequent processes (Zou and Sester, 2022). Specifically, within the context of accurate building-level LiDAR-to-Model registration discussed in this paper, addressing uncertainty inherent in the model serving as reference data will be a central focus.
@@ 3 Methodology
Figure 2 gives an overview of the proposed L2M-Reg. The required inputs are the semantic LoD2 models of the individual building and its corresponding coarsely registered LiDAR point clouds. The LiDAR point clouds could be acquired from MLS systems equipped with GNSS or other well-established coarse registration solutions. The implementation of coarse registration falls outside the scope of this paper, more details can be found from literature (Xu et al., 2017; Bueno et al., 2018; Xu et al., 2019). After preprocessing introduced in Section 3.1, L2M-Reg consists of three key steps: reliable plane correspondence establishment in Section 3.2, pseudo-plane-constrained Gauss–Helmert model in Section 3.3, and adaptable vertical translation estimation in Section 3.4. The core of the L2M-Reg lies in addressing the spatial variation in model uncertainty among different building structural components. L2M-Reg automatically localizes and extracts representative planar regions (i.e., wall plinth) to establish correspondences and obtain 6 DoF transformation parameters. More detailed steps are given in the following subsections.
Figure 2: Flowchart of the proposed L2M-Reg. Each plane of the input semantic LoD2 model is colored for better visualization and the point cloud is colored by intensity.
@@ 3.1 Data Preprocessing
Since the Z-axis of LiDAR point clouds is adjusted to be vertically upward by the sensors, all wall surfaces are assumed to be extruded along their normal vectors in horizontal directions, as shown in Figure 3. The resulting buffer zones with a thickness of 1 m are used to associate individual points with the wall surfaces within whose buffers they reside. The choice of 1 m is mainly based on two considerations: first, many existing LiDAR-to-Model coarse registration methods already achieve decimeter- to centimeter-level accuracy (Diakite and Zlatanova, 2020; Sheik et al., 2022a, b); second, as building plinths are generally within the centimeter- to decimeter-scale, a 1 m buffer reliably covers them without adding processing overhead. Similarly, the DTM is converted into a triangulated irregular network and extruded to filter out redundant LiDAR data. This process also removes elements such as vegetation and pedestrians located near buildings.
Simultaneously, during the buffer zone-based filtering, each individual point is associated with the corresponding wall surface by assigning them unique wall identifiers, as illustrated in Figure 3. Although the finally usable plane segments are not explicitly extracted at this stage, they must reside within the associated LiDAR point clouds. Accordingly, the correspondence relationship are firmly locked at this step.
In contrast to other plane-based methods that first extract plane segments and then search for correspondences (Sheik et al., 2022a; Wujanz et al., 2018; Chen et al., 2019), this strategy maximizes the utilization of existing semantic information from the models. By reliably fixing correspondence relationships before extracting specific plane segments, this strategy ensures a more effective use of prior information while achieving greater simplicity and efficiency as shown in Section 4.4.3.
@@
The main outputs of data preprocessing consist of the model-derived plane $M_{i}$ of the facade area and the corresponding neighboring point cloud $N_{i}$ (equal to the associated LiDAR point clouds shown in Fig 3). The parameters of each $M_{i}$ are explicitly known, and stable correspondence relationships between $N_{i}$ and $M_{i}$ are fixed through data association. In the subsequent step described in Section 3.2 , the finally usable plane correspondences will be established.

### v3

@@ Abstract
Accurate registration between LiDAR (Light Detection and Ranging) point clouds and semantic 3D city models is a fundamental topic in urban digital twinning and a prerequisite for downstream tasks, such as digital construction, change detection, and model refinement. However, achieving accurate LiDAR-to-Model registration at the individual building level remains challenging, particularly due to the generalization uncertainty in semantic 3D city models at the Level of Detail 2 (LoD2). This paper addresses this gap by proposing L2M-Reg, a plane-based fine registration method that explicitly accounts for model uncertainty. L2M-Reg consists of three key steps: establishing reliable plane correspondence, building a pseudo-plane-constrained Gauss-Helmert model, and adaptively estimating vertical translation. Overall, extensive experiments on five real-world datasets demonstrate that L2M-Reg is both more accurate and computationally efficient than current leading ICP-based and plane-based methods. Therefore, L2M-Reg provides a novel building-level solution regarding LiDAR-to-Model registration when model uncertainty is present. The datasets and code for L2M-Reg can be found: https://github.com/Ziyang-Geodesy/L2M-Reg.
@@
{highlights}
A plane-based LiDAR-to-Model fine registration method, L2M-Reg, tailored for individual buildings is proposed.
Explicitly considering the uncertainty in LoD2 models, L2M-Reg achieves superior performance on five real-world datasets.
A 2D-3D decoupled transformation estimation is introduced to mitigate the adverse impact of low-quality ground model data.
A lightweight plane correspondence strategy is developed that leverages the embedded semantic information in LoD2 models.
@@ 1 Introduction
LiDAR (Light Detection and Ranging) point clouds and semantic 3D city models are two widely used digital representations that play important roles in urban digital twinning (Ketzler et al., 2020; Jeddoub et al., 2023). However, as heterogeneous data types, they differ in many aspects such as data format, acquisition methods, geometric features, and accuracy. Specifically, semantic 3D city models primarily contain generalized structural information and serve as simplified and abstract representations of physical buildings, whereas LiDAR point clouds accurately capture detailed geometric features, providing high-precision representations (Kada and McKinley, 2009; Xu and Stilla, 2021). Accurate and efficient registration between them constitutes a fundamental research topic, as well as an essential prerequisite for numerous downstream applications such as digital construction, change detection, model updating and enrichment, model reconstruction, and texturing (Wysocki et al., 2021; Liu et al., 2024; Shao et al., 2024; Zhu et al., 2024; Kulmer et al., 2025).
Internationally, CityGML (City Geography Markup Language), as a standard developed by the Open Geospatial Consortium (OGC), is widely adopted for representing and managing semantic 3D city models (Gröger and Plümer, 2012; Kolbe et al., 2021). CityGML supports the modeling of urban objects by incorporating their 3D geometry, appearance, topology, and semantic information across four distinct Levels of Details (LoD) (Kolbe et al., 2021). Benefiting from the widespread adoption of CityGML, accurately georeferenced LoD2 models are publicly available in many countries and regions (Wysocki et al., 2024). These models are typically maintained by governmental agencies or professional institutions to ensure consistent quality and reliability. Consequently, in most LiDAR-to-Model registration tasks, existing georeferenced LoD2 models are utilized as reference data, to which up-to-date LiDAR point clouds are aligned.
When discussing LoD2 building models used for accurate registration, a particularly critical issue is their inherent model uncertainty. In this context, model uncertainty does not refer to the traditional statistical or probabilistic interpretation, such as variance or covariance describing the positional accuracy of LoD2 model coordinates. The model uncertainty is defined as the inconsistency between the final model, produced through a predefined generation process, and the actual building object, specifically in terms of geometric and structural representation. The encountered uncertainty stems from the current model generation process, which typically relies on 2D building footprints from the cadastral registry. As illustrated in Figure 1, the building footprint is defined by the outermost structural elements. Therefore, it is important to clarify that this study targets structural or generative inconsistencies between LoD2 models and real-world facades rather than statistical error propagation modeled via covariance matrices.
@@
As seen in Figure 1, buildings typically consist of various structural elements from the foundation to the roof, such as wall plinth (in red), exterior facade (in blue), and overhanging roof (in orange). Cadastral surveying regulations define which of these elements should be included in the model. Typically, a horizontal offset exists between a building’s plinth and its upper facade. Depending on the architectural style, this offset can range from several centimeters to decimeters. Since LoD2 building models are generally reconstructed by extruding building footprints to prototypical roof shapes from Airborne Laser Scanning (ALS) data (Roschlaub and Batscheider, 2016), the modeled wall surfaces actually only align with the building’s plinths, not exterior facades. Consequently, the building plinths that closely correspond to the footprint should be more suited for establishing correspondences with modeled wall surfaces.
Previous studies on LiDAR-to-Model registration have commonly assumed that the LoD2 models served as reference data are error-free, thus neglecting this inherent uncertainty. Although this assumption is basically accepted, it will become problematic in high-precision application scenarios. For example, when the captured LiDAR point clouds need to be accurately georeferenced based on existing LoD2 models, this horizontal offset between the plinth and exterior facade cannot be disregarded. Recent research has acknowledged similar issues about model uncertainty, focusing primarily on visual and conceptual representations from architectural or modeling perspectives (Potter et al., 2012; Landes et al., 2019; Zou and Sester, 2022). However, most LiDAR-to-Model registration studies have overlooked this uncertainty issue, treating it as a conventional point cloud registration task. As a result, a disconnection arises between building modeling and LiDAR-to-Model registration research.
Since urban dwellings form the backbone of any city, most existing semantic 3D city models are composed of buildings (Biljecki et al., 2015). Compared to city-level LiDAR-to-Model registration, an accurate and efficient building-level solution offers the potential to refine city models based on specific needs, thereby avoiding extensive data acquisition and substantially reducing costs. However, this transition introduces several challenges:
@@
    Limited availability of geometric features. Individual buildings or partial building segments typically contain fewer geometric primitives, such as facade planes, with limited diversity in orientation and quantity. This constraint makes feature extraction and correspondence establishment more difficult.
@@
These challenges underscore a clear research gap in achieving accurate building-level LiDAR-to-Model registration under model uncertainty. To address this, a tailored solution named L2M-Reg is proposed. To the best of our knowledge, this is the first comprehensive study to explicitly account for the inherent uncertainty of reference LoD2 models in building-level registration. The main contributions are summarized as follows:
@@
    A plane-based LiDAR-to-Model fine registration method, L2M-Reg, tailored for individual buildings is proposed. It integrates reliable plane correspondence establishment, pseudo-plane-constraint Gauss-Helmert Model (GHM), and adaptable vertical translation estimation. By explicitly considering the uncertainty in LoD2 models, L2M-Reg achieves superior performance on all five real-world datasets.
@@ 2.1.1 Registration Directly Using LoD2 Models
The registration between LiDAR point clouds and their corresponding LoD2 models essentially involves feature extraction and correspondence matching across heterogeneous data. A straightforward strategy to address this issue is to directly extract geometric features, such as points, lines, and planes, from the models and subsequently match these features with corresponding ones derived from the point clouds (Bosché, 2012; Bueno et al., 2018; Goebbels et al., 2019; Lucks et al., 2021; Sheik et al., 2022a, b; Monasse et al., 2023). The transformation parameters can then be computed based on these matched features. In particular, most man-made structures include numerous planar surfaces, which naturally facilitates the extraction and utilization of planar features (Sheik et al., 2022a; Zhao et al., 2022; Qiao and Butt, 2023; Yang, 2023; Zhang et al., 2024).
@@ 2.1.2 Registration Using Converted Point Clouds from LoD2 Models
ICP-based methods are extensively employed to register point clouds. Traditional ICP algorithms iteratively align corresponding points but often exhibit slow convergence and sensitivity to initial values and outliers (Besl and McKay, 1992). Thus, many enhanced variants have been developed to address these limitations. For instance, Point-to-Line ICP (Censi, 2008) and Point-to-Plane ICP (Rusinkiewicz and Levoy, 2001; Low, 2004) utilize local linear and planar features, respectively, thereby accelerating convergence and enhancing registration accuracy. Generalized ICP (GICP) further refines registration precision by statistically optimizing correspondences through probabilistic modeling of local surface geometry (Segal et al., 2009). Trimmed ICP (TriICP) effectively reduces the influence of outliers by selectively excluding a certain percentage of correspondences, thereby enhancing robustness in noisy scenarios (Chetverikov et al., 2002). More recently, a robust method based on generated planar patches and adaptive distance thresholds was proposed, which has significantly reduced the influence of surface changes on registration accuracy (Yang and Schwieger, 2023; Yang and Holst, 2025). KISS-ICP is also widely used for point cloud registration because of its versatility and computational efficiency, particularly in robotics and SLAM applications (Vizzo et al., 2023). Collectively, these ICP variants have progressively addressed the limitations in standard ICP, significantly improving the accuracy, efficiency, and practical applicability of point cloud registration tasks.
Compared with the aforementioned ICP-based methods, geometric feature-based methods utilize points, lines, curves, and planes extracted from point clouds to establish correspondences for transformation estimation. These methods are generally less sensitive to variations of initial alignments compared to ICP-based methods. In the building scenarios discussed in this paper, planar features offer notable advantages over other geometric primitives, such as point or line features, and have thus been commonly utilized in point cloud registration tasks (Bosché, 2012; Wujanz et al., 2018; Chen et al., 2019; Sheik et al., 2022a; Xu et al., 2025a). This preference arises primarily for two reasons. First, man-made structures typically contain abundant planar elements, naturally facilitating the extraction and use of planar features. Second, planar features demonstrate higher resilience against surface noise and outliers.
@@
According to the summary above, the majority of published studies still regard ICP-based methods and PLADE as appropriate state-of-the-art baselines for comparison in the context of LiDAR-to-Model registration.
@@ 2.2 Inherent Uncertainty of Models
Compared to the rapid growth in the creation and application of models, research on their uncertainty has lagged behind and not received equivalent attention (Zou and Sester, 2022). Initially, researchers from architecture and archaeology took the lead in visualizing uncertainty associated with models of historical buildings (Zuk et al., 2005; Kensek, 2007). Subsequently, solutions to quantifying uncertainty in models have continued to evolve, with increasingly comprehensive methods proposed, along with the introduction of the concept of Level of Uncertainty (LoU), which systematically characterizes uncertainty variations among different components of building models (Potter et al., 2012; Landes et al., 2019).
With growing demand for high-precision applications relying on semantic 3D city models, such as navigation (Kulmer et al., 2025), trajectory estimation (Lucks et al., 2021), and change detection (Meyer et al., 2022), an increasing number of researchers have also begun to investigate the uncertainty inherent in these models (Foschi et al., 2024). For instance, probabilistic methods have been developed for quantifying uncertainty, primarily focusing on semantic labeling in 2D maps or building models (Paz et al., 2020; Feng et al., 2021; Di et al., 2022). Additionally, other researchers have specifically examined the uncertainty associated with facades and windows in semantic 3D city models, quantifying aspects such as their positions and orientations (Zou and Sester, 2022).
Collectively, uncertainty in semantic 3D city models has become increasingly important in high-precision applications, as uncertainty inherent in the model inevitably propagates into downstream use cases. Neglecting these uncertainties or treating the model as an error-free reference is likely to introduce systematic inaccuracies and errors in subsequent processes (Zou and Sester, 2022). Specifically, within the context of accurate building-level LiDAR-to-Model registration discussed in this study, addressing uncertainty inherent in the model serving as reference data will be a central focus.
@@ 3 Methodology
Figure 2 gives an overview of the proposed L2M-Reg. The required inputs are the semantic LoD2 models of the individual building and its corresponding coarsely registered LiDAR point clouds. The LiDAR point clouds could be acquired from MLS systems equipped with GNSS or other well-established coarse registration solutions. The implementation of coarse registration falls outside the scope of this study, and more details can be found from the literature (Xu et al., 2017; Bueno et al., 2018; Xu et al., 2019). After preprocessing introduced in Section 3.1, L2M-Reg consists of three key steps: reliable plane correspondence establishment in Section 3.2, pseudo-plane-constrained Gauss-Helmert model in Section 3.3, and adaptable vertical translation estimation in Section 3.4. The core of the L2M-Reg lies in addressing the spatial variation in model uncertainty among different building structural components. L2M-Reg automatically localizes and extracts representative planar regions (i.e., wall plinth) to establish correspondences and obtain 6 DoF transformation parameters. More detailed steps are given in the following subsections.
Figure 2: Flowchart of the proposed L2M-Reg. Each plane of the input semantic LoD2 model is colored for better visualization, and the point cloud is colored by intensity.
@@ 3.1 Data Preprocessing
Since the Z-axis of LiDAR point clouds is adjusted to be vertically upward by the sensors, all wall surfaces are assumed to be extruded along their normal vectors in horizontal directions, as shown in Figure 3. The resulting buffer zones with a thickness of 1 m are used to associate individual points with the wall surfaces within whose buffers they reside. The choice of 1 m is mainly based on two considerations: first, many existing LiDAR-to-Model coarse registration methods already achieve decimeter-to-centimeter-level accuracy (Diakite and Zlatanova, 2020; Sheik et al., 2022a, b); second, as building plinths are generally within the centimeter-to-decimeter scale, a 1 m buffer reliably covers them without adding processing overhead. Similarly, the DTM is converted into a triangulated irregular network and extruded to filter out redundant LiDAR data. This process also removes elements such as vegetation and pedestrians located near buildings.
Simultaneously, during the buffer zone-based filtering, each individual point is associated with the corresponding wall surface by assigning them unique wall identifiers, as illustrated in Figure 3. Although the finally usable plane segments are not explicitly extracted at this stage, they must reside within the associated LiDAR point clouds. Accordingly, the correspondence relationships are firmly locked at this step.
The built-in semantics in LoD2 models play a critical role in this step, particularly in the accurate identification of wall surfaces. These semantics allow wall surfaces to be reliably selected and associated with the corresponding LiDAR point clouds through buffer zone-based filtering. In this way, semantic information is transferred from the model to the point clouds, enabling identification of both the wall surfaces in the model and their corresponding point cloud regions. This process provides the foundation for the subsequent extraction of wall plinths and enables more reliable identification of wall-related regions without introducing additional data labeling costs.
In contrast to other plane-based methods that first extract plane segments and then search for correspondences (Wujanz et al., 2018; Chen et al., 2019; Sheik et al., 2022a), this strategy maximizes the utilization of existing semantic information from the models. By reliably fixing correspondence relationships before extracting specific plane segments, this strategy ensures a more effective use of prior information while achieving greater simplicity and efficiency as shown in Section 4.4.3.
@@
The main outputs of data preprocessing consist of the model-derived plane $M_{i}$ of the facade area and the corresponding neighboring point cloud $N_{i}$ (equal to the associated LiDAR point clouds shown in Figure 3). The parameters of each $M_{i}$ are explicitly known, and stable correspondence relationships between $N_{i}$ and $M_{i}$ are fixed through data association. In the subsequent step described in Section 3.2, the finally usable plane correspondences will be established.